# On an entropy inequality for quadratic forms and applications

### v2

@@ 4.2 Diagonal forms with dependent random variables
(ii) For all $n\in\mathbb{N}$,
$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,H_{n}(XY)\right\}\geq n(s+\epsilon_{3})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{s}(1).$
@@
(i) Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.22 it is also jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman. Applying Theorem 1.4’, there is some $\epsilon_{2}=\epsilon_{2}(s)$ such that
@@
(ii) Notice that
$xy=\frac{1}{4}(x+y)^{2}-\frac{1}{4}(x-y)^{2}.$
Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.23 it is also jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, with $\log C^{\prime\prime}=\log C_{1}C_{2}+O(1)$.
Applying Theorem 1.4’ with $\alpha=1/4$, $\beta=-1/4$ and the random variable (X+Y,XY)(X+Y,X-Y), we obtain
$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,H_{n}(2X),\,H_{n}(XY)\right\}$ (52)
$\qquad\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$
for some $\epsilon_{2}=\epsilon_{2}(s)$. If we switch the role of $X$ and $Y$, then similar conclusion holds, with $H_{n}(2X)$ replaced by $H_{n}(2Y)$. Therefore, the left-hand side of (52) can be replaced by
$\max\left\{\frac{H_{n}(X+Y)+H_{n}(X-Y)}{2},\,\frac{H_{n}(2X)+H_{n}(2Y)}{2},\,H_{n}(XY)\right\}.$
By Lemma 2.24, combining $s_{1}+s_{2}\geq 2s$, the second item satisfies
$H_{n}(X,Y)\geq\frac{H_{n}(2X)+H_{n}(2Y)+O(1)}{2}+ns+O(\log_{+}C_{1}C_{2})+O_{s}(1).$
Combining the fact
$H_{n}(X,Y)=H_{n}(X+Y,X-Y)+O(1)\leq H_{n}(X+Y)+H_{n}(X-Y)+O(1),$
the first conclusion follows by taking $\epsilon_{3}=\epsilon_{2}/2$. ∎
4.3 The general forms
Now we are ready to prove Theorem 1.2. It actually holds for any non-degenerate real quadratic form $\phi(x,y)=a_{1}x^{2}+a_{2}xy+a_{3}y^{2}$ that is congruent over $\mathbb{Q}$ to a non-degenerate real diagonal form, i.e.,
for some $\alpha,\beta\in{\mathbb{R}}^{\ast}$ and $\xi_{11},\xi_{12},\xi_{21},\xi_{22}\in\mathbb{Q}$ with $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$. Here, all the constants $\alpha,\beta,\xi_{11},\xi_{12},\xi_{21},\xi_{22}$ depends on the coefficients $a_{1},a_{2},a_{3}$ in $\phi(x,y)$.
Theorem 1.2’ Let $\phi(x,y)$ be in the form (53). Let $0<s<1$, $C\geq 1$ and $c>0$. There is a positive constant $\epsilon_{4}=\epsilon_{4}(\phi,s)$ such that the following statement holds: Let $X$ and $Y$ be independent real random variables with ranges contained in [c,c][-c,c]. Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(x,y))\right\}\geq n(s+\epsilon)+O_{s}(\log_{+}C)+O(\log_{+}c)+O_{\phi,s}(1)$
@@ 4.3 The general forms
By Lemma 2.23, the random variable $Z:=(\xi_{11}X+\xi_{12}Y,\xi_{21}X+\xi_{22}Y)$ is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, where $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$. Moreover, the range of $Z$ is contained in $c^{\prime\prime},c^{\prime\prime}]^{2}$ with $c^{\prime\prime}=O_{\phi}(c)$. Applying Theorem 1.4’ with $Z$, we obtain some positive number $\epsilon_{2}=\epsilon_{2}(s)$ such that
$\max\left\{\frac{H_{n}(\xi_{11}X+\xi_{12}Y)+H_{n}(\xi_{21}X+\xi_{22}Y)}{2},\,H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)},\,H_{n}\big{(}\phi(X,Y)\big{)}\right\}$
$\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$ (54)
@@
Next, using Lemma 2.21, let us assume that $d=d(\xi_{11},\xi_{12},\xi_{21},\xi_{22})=d(\phi)$ is the positive number satisfying
$H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)}\leq dH_{n}(X+Y)-(d-1)\frac{H_{n}(X)+H_{n}(Y)}{2}+O(1).$ (55)
Combining Lemma 2.24(i) and the fact that $s_{1}+s_{2}\geq 2s$, we see that the right-hand side of (55) is
$\leq dH_{n}(X+Y)-(d-1)\frac{(ns_{1}-\log C_{1}+O(1))+(ns_{2}-\log C_{2}+O(1))}{2}+O(1)$
$\leq dH_{n}(X+Y)-(d-1)ns+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1).$
This leads to
$H_{n}(X+Y)\geq\frac{H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)}+(d-1)ns}{d}+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1).$
Similarly, we can deduce that
$H_{n}(X+Y)\geq\frac{H_{n}(\xi_{11}X+\xi_{12}Y)+(d^{\prime}-1)ns}{d^{\prime}}+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1),$
and also
$H_{n}(X+Y)\geq\frac{H_{n}(\xi_{21}X+\xi_{22}Y)+(d^{\prime\prime}-1)ns}{d^{\prime\prime}}+O_{\phi}(\log_{+}C_{1}C_{2})+O_{\phi}(1)$
for some positive numbers $d^{\prime}=d^{\prime}(\phi)$ and $d^{\prime\prime}=d^{\prime\prime}(\phi)$. Now we deduce from (54) that
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{5})+O_{\phi,s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$
for $\epsilon_{4}=\epsilon_{2}/(d+d^{\prime}+d^{\prime\prime})$ and all $n\in\mathbb{N}$. Theorem 1.2 then follows. ∎
@@
In the conclusions of Lemma 4.1 and Theorems 1.2’-1.4’, the implied constants in the terms $O_{s}(\log_{+}C_{1}C_{2})$, $O_{\phi,s}(\log_{+}C_{1}C_{2})$ or $O_{s}(\log_{+}C)$ can be independent of $s$ if either of the following two situations happens: (1) we are only considering the case $1/2<s<1$; (2) the range of random variables are bounded away from 0 by some absolute constant. We preserve the dependence on $s$ in these terms for the uniformity with the case $0<s\leq 1/2$.
@@ 5.2.2 Modifications to their argument
Since $H^{s}(A)>0$, by the Frostman lemma [13, Theorem 2.7], there exists a probability $(s,1)$-Frostman measure $\mu$ supported on $A$. Let $X$ and $Y$ be i.i.d. random variables that both have distribution measure $\mu$.
Theorem 1.2 gives us
$\max\big{\{}H_{n}(X+Y),\,H_{n}(\phi(X,Y))\big{\}}\geq(1+\epsilon)ns$
for sufficiently large $n$.
By Lemma 2.4(ii), we deduce that
$\log\#\{D_{n}(x)\mid x\in A+A\}\geq H_{n}(X+Y),$
and
$\log\#\{D_{n}(x)\mid x\in\phi(A,A)\}\geq H_{n}(\phi(X,Y)).$
Recall that $\mathcal{E}_{2^{-n}}(A+A)$ and $\mathcal{E}_{2^{-n}}(\phi(A,A))$ are the least numbers of intervals of length $2^{-n}$ needed to cover $A+A$ and $\phi(A,A)$, respectively. It is clear that
$\log\mathcal{E}_{2^{-n}}(A+A)\geq H_{n}(X+Y),~~\log\mathcal{E}_{2^{-n}}(\phi(A,A))\geq H_{n}(\phi(X,Y)).$
Hence,
$\max\left\{\frac{\log\mathcal{E}_{2^{-n}}(A+A)}{n},\frac{\log\mathcal{E}_{2^{-n}}(\phi(A,A))}{n}\right\}\geq\frac{1}{n}\max\left\{H_{n}(X+Y),H_{n}(\phi(X,Y))\right\}\geq(1+\epsilon)s.$
When $n\to\infty$, we obtain
$\max\{\underline{\dim}_{B}(A+A),\,\overline{\dim}_{B}(\phi(A,A))\}\geq(1+\epsilon)s,$
and
$\max\{\overline{\dim}_{B}(A+A),\,\underline{\dim}_{B}(\phi(A,A))\}\geq(1+\epsilon)s.$
This completes the proof. ∎
Proof of Theorem 1.6.
We still choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ satisfying $A_{\eta\delta}+A_{\eta\delta}$ and $\phi(A_{\eta\delta},A_{\eta\delta})$ are contained in the $(A+A)_{\delta/10}$ and $(\phi(A,A))_{\delta/10}$, respectively, for any $\delta>0$. Since $A$ is $\delta$-separated, the Lebesgue measure of $A_{\eta\delta}$ is $\gg_{\phi}\delta^{1-s}$. Let $X$ be a random variable that has uniform distribution on $A_{\eta\delta}$. Combining the non-concentration condition, one deduces that
$P(X\in B(x,r))\ll_{\phi}\begin{cases}r\delta^{s-1},\quad&\text{if }0<r\leq\delta,\\ Cr^{s},\quad&\text{if }\delta<r\leq 1,\\ 1,\quad&\text{if }r>1,\end{cases}$
for any ball $B(x,r)$. In all the cases, the right-hand side can be bounded by $Cr^{s}$. So $X$ is $(s,O_{\phi}(C))$-Frostman.
Let $Y$ be a random variable that is i.i.d. with $X$. Now similar relations as in (58), (61) and (62) hold. As a replacement of (60), we apply Theorem 1.2 to get an $\epsilon^{\prime}=\epsilon^{\prime}(s,\phi)$ such that
$\max\{H_{n}(X+Y),\,H_{n}(\phi(x,y))\}\geq n(s+\epsilon^{\prime})+O_{\phi,s}(\log C)+O_{\phi,s}(1).$
for all $n\in\mathbb{N}$. Putting everything together, we arrive at the conclusion that
$\max\{\mathcal{E}_{\delta}(A+A),\,\mathcal{E}_{\delta}(\phi(A,A))\}\gg_{\phi,s}C^{-O_{\phi,s}(1)}\delta^{-\epsilon}(\#A)$
for all $\delta>0$. ∎
Proof of Theorem 1.7.
Let $\epsilon=\epsilon(s)$ to be determined later.
Note that $f_{1}(x,y)=x+y$ and $f_{2}(x,y)=\phi(x,y)=\alpha x^{2}+\beta y^{2}$ are $O_{\phi}(1)$-Lipchitz on $0,1]^{2}$. We can choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ such that $A_{\eta\delta}+_{G}A_{\eta\delta}$ and $\alpha\cdot A_{\eta\delta}^{\wedge 2}+\beta\cdot A_{\eta\delta}^{\wedge 2}$ are contained in the $(A+_{G}A)_{\delta/10}$ and $(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})_{\delta/10}$, respectively, for any $\delta>0$. Then
$\begin{split}\mathcal{E}_{\delta}(A+_{G}A)&\geq\frac{1}{2}\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta}),\\ \mathcal{E}_{\delta}(\alpha\cdot A^{\wedge 2}+_{G}\beta\cdot A^{\wedge 2})&\geq\frac{1}{2}\mathcal{E}_{\delta}(\alpha\cdot A_{\eta\delta}^{\wedge 2}+_{G}\beta\cdot A_{\eta\delta}^{\wedge 2}).\end{split}$ (58)
@@
Now consider the set $G_{\eta\delta}\subseteq A_{\eta\delta}\times A_{\eta\delta}$. Since $A$ is $\delta$-separated, and $|G|\geq\delta^{\epsilon}|A|^{2}\geq\delta^{-2s+\epsilon}$, the $2$-dimensional Lebesgue measure of $G_{\eta\delta}$ is $\gg\eta^{2}\delta^{2-2s+\epsilon}\gg_{\phi}\delta^{2-2s+\epsilon}$.
Let $(X,Y)$ be a random variable that has uniform distribution on $G_{\eta\delta}$. Combining the non-concentration condition, one can verify that
[…]
for any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$. In all the cases, the right-hand side can be bounded by $C^{2}\delta^{-\epsilon}r_{1}^{s}r_{2}^{s}$. So $(X,Y)$ is jointly $(s,s,C^{2}\delta^{-\epsilon})$-Frostman.
By Theorem 1.4, there exists an $\epsilon^{\prime}=\epsilon^{\prime}(s)$ such that
$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\alpha X^{2}+\beta Y^{2})\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{s}(\log_{+}C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1).\end{split}$ (60)
@@
$\begin{split}&\quad\mathcal{E}_{\delta}(A_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(A_{\eta\delta})\geq\frac{1}{2}\max\left\{\#\,range_{+}(D_{n}(X)),\,\#\,range_{+}(D_{n}(Y))\right\},\\ &\qquad\;\;\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(A_{\eta\delta}+_{G}A_{\eta\delta})\geq\frac{1}{2}\#\,range_{+}(D_{n}(X+Y)),\\ &\mathcal{E}_{\delta}(\alpha\cdot A_{\eta\delta}^{\wedge 2}+_{G}\beta\cdot A_{\eta\delta}^{\wedge 2})\geq\mathcal{E}_{2^{-n}}(\alpha\cdot A_{\eta\delta}^{\wedge 2}+_{G}\beta\cdot A_{\eta\delta}^{\wedge 2})\geq\frac{1}{2}\#\,range_{+}(\alpha X^{2}+\beta Y^{2}).\end{split}$ (62)
Putting (58)-(62) together, one arrives at
$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\alpha\cdot A^{2}+_{G}\beta\cdot A^{2})\}$
@@
$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\alpha\cdot A^{2}+_{G}\beta\cdot A^{2})\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}$
@@
$\max\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\alpha\cdot A^{2}+_{G}\beta\cdot A^{2})\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A)$

### v3

@@ 4.3 The general forms
for all $n\in\mathbb{N}$.
@@
Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.30 it is also jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman. Applying Theorem 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(s)$ such that
@@
$\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$.
Theorem 1.3”. Suppose that $\phi(x,y)$ has the form (56), where $\alpha,\beta\in{\mathbb{R}}^{\ast}$, $\xi_{11},\xi_{12},\xi_{13},\xi_{14}\in\mathbb{Q}^{\ast}$, $\xi_{11}\neq-\xi_{21}$, $\xi_{12}\neq-\xi_{22}$ and $\xi_{11}\xi_{22}-\xi_{12}\xi_{21}\neq 0$. Let $0<s<1$ and $C_{1},C_{2},c>0$. There is a positive constants $\epsilon_{4}=\epsilon_{4}(\phi,s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1).$
@@
Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, Lemma 2.30 tells that it is both marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman and jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman. It follows from Lemma 2.32(i) and the fact $s_{1}+s_{2}\geq 2s$ that
$\min\big{\{}H_{n}(X)+H_{n}(Y),\,H_{n}(X,Y)\big{\}}\geq 2ns-\log C_{1}C_{2}+O(1).$ (57)
In view of Lemma 2.32(ii), one has
$\frac{H_{n}(X)+H_{n}(Y)}{2}\leq H_{n}(X,Y)-ns+O(\log_{+}C_{1}C_{2})+O(1).$
It follows that
$H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\leq H_{n}(X,Y)-2ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).$ (58)
Consider the random variable $(\xi_{11}X+\xi_{12}Y,\,\xi_{21}X+\xi_{22}Y)$. Its range is contained in $c^{\prime\prime},c^{\prime\prime}]^{2}$, with $c^{\prime\prime}=O_{\phi}(c)$. And Lemma 2.31 shows that it is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, with $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$, since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. By Lemma 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(s)>0$ such that
$\max\left\{\frac{H_{n}(\xi_{11}X+\xi_{12}Y)+H_{n}(\xi_{21}X+\xi_{22}Y)}{2},\,H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)},\,H_{n}(\phi(X,Y))\right\}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\geq n(s+\epsilon_{2})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$ (59)
@@
Notice that $\xi_{11}\neq-\xi_{21}$ and $\xi_{12}\neq-\xi_{22}$. It follows from Theorem 2.23 that
$H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)}-\frac{H_{n}(X)+H_{n}(Y)}{2}$
$\qquad\leq d_{1}\left(H_{n}\big{(}X+Y\big{)}-\frac{H_{n}(X)+H_{n}(Y)}{2}\right)+d_{2}\big{(}H_{n}(X)+H_{n}(Y)-H_{n}(X,Y)\big{)}+O_{\phi}(1)$ (60)
for some $d_{1}=d_{1}(\phi)>0$ and $d_{2}=d_{2}(\phi)\geq 0$. By (57), (58) and (60), one gets
$\begin{split}&d_{1}\cdot H_{n}(X+Y)+2d_{2}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}\big{(}(\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big{)}\\ &\qquad\qquad\qquad+(d_{1}+2d_{2}-1)ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).\end{split}$ (61)
Similar arguments show that
$\begin{split}&d_{1}^{\prime}\cdot H_{n}(X+Y)+2d_{2}^{\prime}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}(\xi_{11}X+\xi_{12}Y)\\ &\qquad\qquad\qquad+(d_{1}^{\prime}+2d_{2}^{\prime}-1)ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).\end{split}$ (62)
for some $d_{1}^{\prime}=d_{1}^{\prime}(\phi)>0$ and $d_{2}^{\prime}=d_{2}^{\prime}(\phi)\geq 0$. And also
$\begin{split}&d_{1}^{\prime\prime}\cdot H_{n}(X+Y)+2d_{2}^{\prime\prime}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}(\xi_{21}X+\xi_{22}Y)\\ &\qquad\qquad\qquad+(d_{1}^{\prime\prime}+2d_{2}^{\prime\prime}-1)ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).\end{split}$ (63)
for some $d_{1}^{\prime\prime}=d_{1}^{\prime\prime}(\phi)>0$ and $d_{2}^{\prime\prime}=d_{2}^{\prime\prime}(\phi)\geq 0$.
Now the theorem follows by combining (59) and (61)-(63), with $\epsilon_{4}=\epsilon_{2}/\max\{d_{1}+2d_{2},d_{1}^{\prime}+2d_{2}^{\prime},d_{1}^{\prime\prime}+2d_{2}^{\prime\prime}\}$. ∎
In the following, Theorem 1.4 will be proved in the following form.
Theorem 1.4’ Let $\phi(x,y)$ be in the form (56). Let $0<s<1$, $C_{1},C_{2},c>0$. There is a positive constant $\epsilon_{4}=\epsilon_{4}(\phi,s)$ such that the following statement holds: Let $X$ and $Y$ be independent real random variables with ranges contained in [c,c][-c,c]. Suppose that $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, where $0\leq s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\max\left\{H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1)$
for all $n\in\mathbb{N}$.
Proof.
When $\phi(x,y)$ is diagonal, the conclusion can be deduced from Lemma 4.1. In the following, we concentrate on the non-diagonal case. By Lemma 2.30, the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. Applying Theorem 1.3”, there is some positive constant $\epsilon_{4}=\epsilon_{4}(\phi,s)$ such that
$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1).$
Combining (13), one has
$H_{n}(X,Y)\leq H_{n}(X+Y)+H_{n}(X)+O(1)\leq 2H_{n}(X+Y)+O(1).$
The theorem then follows. ∎
@@
In the conclusions of Theorems 1.2’-1.4’, the implied constants in the terms $O_{s}(\log_{+}C)$ or $O_{s}(\log_{+}C_{1}C_{2})$ can be independent of $s$ if either of the following two situations happens: (1) we are only considering the case $s\gg 1$; (2) the range of random variables are bounded away from 0 by some absolute constant. We preserve the dependence on $s$ in these terms for the uniformity with the case $0<s\leq 1/2$.
@@ 5.2.2 Modifications to their argument
Write $\phi(x,y)=\alpha x^{2}+\beta y^{2}$. Let $\epsilon=\epsilon(s)$ to be determined later.
Note that $(x,y)\mapsto x+y$ and $(x,y)\mapsto\phi(x,y)$ are both $O_{\phi}(1)$-Lipchitz on $0,1]^{2}$. We can choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ such that $A_{\eta\delta}+_{G}A_{\eta\delta}$ and $\phi_{G}(A_{\eta\delta},A_{\eta\delta})$ are contained in the $(A+_{G}A)_{\delta/10}$ and $(\phi_{G}(A,A))_{\delta/10}$, respectively, for any $\delta>0$. Then
$\begin{split}\mathcal{E}_{\delta}(A+_{G}A)&\gg\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta}),\\ \mathcal{E}_{\delta}(\phi_{G}(A,A))&\gg\mathcal{E}_{\delta}(\phi_{G}(A_{\eta\delta},A_{\eta\delta}))).\end{split}$ (66)
@@
Now consider the set $G_{\eta\delta}\subseteq A_{\eta\delta}\times A_{\eta\delta}$. Since $A$ is $\delta$-separated, and $|G|\geq\delta^{\epsilon}|A|^{2}\geq\delta^{-2s+\epsilon}$, then
$\mathcal{L}^{2}(G_{\eta\delta})\gg\eta^{2}\delta^{2-2s+\epsilon}\gg_{\phi}\delta^{2-2s+\epsilon}.$ (68)
Let $(X,Y)$ be a random variable that has uniform distribution on $G_{\eta\delta}$. For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, we have
$P(X\in B(x,r_{1}),Y\in B(y,r_{2}))\ll\frac{\mathcal{L}^{1}(B(x,r_{1})\cap A_{\eta\delta})\cdot\mathcal{L}^{1}(B(y,r_{2})\cap A_{\eta\delta})}{\mathcal{L}^{2}(G_{\eta\delta})}.$
When $\delta<r_{2}\leq 1$, the concentration conditions gives
$\mathcal{L}^{1}(B(y,r_{2})\cap A_{\eta\delta})\ll Cr_{2}^{s}(\#A)\delta\ll cr_{2}^{s}\delta^{1-s}.$
Calculation reveals that
$P(X\in B(x,r_{1}),Y\in B(y,r_{2}))\ll_{\phi}\begin{cases}r_{1}\cdot r_{2}\cdot\delta^{2s-2-\epsilon},\quad&\text{if }0<r_{1},r_{2}\leq\delta,\\ r_{1}\cdot Cr_{2}^{s}\delta^{1-s}\cdot\delta^{2s-2-\epsilon}&\text{if }0<r_{1}\leq\delta<r_{2}\leq 1,\\ r_{1}\cdot\delta^{1-s}\cdot\delta^{2s-2-\epsilon},\quad&\text{if }0<r_{1}\leq\delta,\,r_{2}>1,\\ Cr_{1}^{s}\delta^{1-s}\cdot Cr_{2}^{s}\delta^{1-s}\cdot\delta^{2s-2-\epsilon}&\text{if }\delta<r_{1},r_{2}\leq 1\\ Cr_{1}^{s}\delta^{1-s}\cdot\delta^{1-s}\cdot\delta^{2s-2-\epsilon},\quad&\text{if }\delta<r_{1}\leq 1<r_{2},\\ 1,\quad&\text{if }r_{1},r_{2}>1,\end{cases}$
and the upper bounds in other cases can be obtained by symmetry. In all the cases, the right-hand side can be bounded by $C^{2}\delta^{-\epsilon}r_{1}^{s}r_{2}^{s}$. So $(X,Y)$ is jointly $(s,s,O_{\phi}(C^{2}\delta^{-\epsilon}))$-Frostman.
By Theorem 1.2’, there exists an $\epsilon^{\prime}=\epsilon^{\prime}(s)$ such that
$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{s}(\log_{+}C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1).\end{split}$ (69)
@@
$\begin{split}&\quad\mathcal{E}_{\delta}(A_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(A_{\eta\delta})\gg\max\left\{\#\,range_{+}(D_{n}(X)),\,\#\,range_{+}(D_{n}(Y))\right\},\\ &\qquad\;\;\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta})\geq\mathcal{E}_{2^{-n}}(A_{\eta\delta}+_{G}A_{\eta\delta})\gg\#\,range_{+}(D_{n}(X+Y)),\\ &\qquad\quad\mathcal{E}_{\delta}(\phi_{G}(A_{\eta\delta},A_{\eta\delta}))\geq\mathcal{E}_{2^{-n}}(\phi_{G}(A_{\eta\delta},A_{\eta\delta}))\gg\#\,range_{+}(\phi(X,Y)).\end{split}$ (71)
Putting (66), (67), and (69)- (71) together, one arrives at
$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}$
@@
$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}$ (73)
@@
$\max\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A)$